OpenStudy (anonymous):

factor 16t^2+64t-60 please explain steps

OpenStudy (mertsj):

First factor out 4

OpenStudy (anonymous):

4t^2-16t+15

OpenStudy (mertsj):

Either you posted the problem wrong or you have factored it wrong.

OpenStudy (mertsj):

Check the signs in your original post.

OpenStudy (anonymous):

i brought it to the oppsite side makin all the signs opposite

OpenStudy (mertsj):

There is no equal sign so there is no opposite side. Are the signs in your original post correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

=0

OpenStudy (mertsj):

\[16t^2+64t-60=0\]

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

\[4(4t^2+16t-15)=0\]

OpenStudy (anonymous):

where do we go from there

OpenStudy (mertsj):

It will not factor any further. That's why I suspect you have the signs copied wrong. It it were 4t^2-16t+15 it would factor again. So if you have it copied right, you will have to use the quadratic formula. So why don't you check your signs again because I am fairly confident that you have posted it incorrectly.

OpenStudy (anonymous):

heres a similar problem 10m^2-23m+12=0

OpenStudy (mertsj):

Factor it.

OpenStudy (mertsj):

(10m-3)(m-2)=0

OpenStudy (anonymous):

what happens to the 23? how does that get factored

OpenStudy (mertsj):

Have you heard of FOIL?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

To factor a trinomial, you use FOIL in reverse. Why are you studying quadratic equations if you haven't had factoring yet?

OpenStudy (anonymous):

i just forgot is all

OpenStudy (mertsj):

Well, use FOIL on the two binomials I gave you and see if they multiply back to the trinomial you are trying to factor. Also you need to brush up on your factoring because you need to master it before going on.

OpenStudy (anonymous):

can i just solve this using the quadratic formula

OpenStudy (mertsj):

All quadratic equations can be solved using the quadratic formula if you put them in the ax^2+bx+c = 0 form

OpenStudy (anonymous):

so i dont even need to factor this i can just use the quadratic formula right

OpenStudy (mertsj):

Depends on what your teacher said for you to do. If he/she or the textbook said solve by factoring, then that is what you have to do. If it just says "solve" then you may select the method you choose but I am telling you that going forward in mathematics, you will need to be able to factor.

OpenStudy (anonymous):

how do i make decimals into fractions like 1.5

OpenStudy (mertsj):

1.5 means 1 and 5/10. reduce the 5/10 and you have 1/2 so 1.5= 1 1/2

OpenStudy (anonymous):

sorry im slow

OpenStudy (anonymous):

3/2 = 1.5

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

how do i turn a 1.5 into a 3/2

OpenStudy (anonymous):

nvm i got it wow im a retard thank u so much

OpenStudy (mertsj):